OpenStudy (yacoub1993):

Your computer supply store sells two types of laser printers. The first type, A, has a cost of $86 and you make a $45 profit on each one. The second type, B, has a cost of $130 and you make a $35 profit on each one. You expect to sell at least 100 laser printers this month and you need to make at least $3850 profit on them. How many of what type of printer should you order if you want to minimize your cost?

OpenStudy (yacoub1993):

so far i did this much but i am stuck x = number of type A printers y = number of type B printers x + y ≥ 100 Type A, you make a profit of $45 on each Type B, you make a profit of $35 on each Therefore you need to make at least $3850 profit Therefore 45x + 35y ≥ 3850 Minimize Cost = 86x + 130y x + y ≥ 100 45x + 30y ≥ 3450

OpenStudy (yacoub1993):

@apple_pi Help Please

OpenStudy (anonymous):

Hang on a sec... jotting things down

OpenStudy (yacoub1993):

ok

OpenStudy (anonymous):

Actually gotta go now. I'll be back if I can

OpenStudy (yacoub1993):

@robtobey Help me Please

OpenStudy (yacoub1993):

@INeedHelpPlease? Help Me Here

terenzreignz (terenzreignz):

Well, this is quite a sticky situation you got yourself into (optimisation... the horror D: ) But anyway, you see those inequalities called constraints? They both correspond to a specific line, right? Find their intersection...

terenzreignz (terenzreignz):

\[\Large x + y = 100\\\Large 45x + 35y = 3850\] Solve this system :)

terenzreignz (terenzreignz):

Still here, are you? First, solve that system I put up... ^